### v1

# Generalizing in Net-Zero Microgrids: A Study with Federated PPO and TRPO

Abstract
This work addresses the challenge of optimal energy management in microgrids through a collaborative and privacy-preserving framework. We propose the FedTRPO methodology, which integrates Federated Learning (FL) and Trust Region Policy Optimization (TRPO) to manage distributed energy resources (DERs) efficiently. Using a customized version of the CityLearn environment and synthetically generated data, we simulate designed net-zero energy scenarios for microgrids composed of multiple buildings. Our approach emphasizes reducing energy costs and carbon emissions while ensuring privacy. Experimental results demonstrate that FedTRPO is comparable with state-of-the-art federated RL methodologies without hyperparameter tunning. The proposed framework highlights the feasibility of collaborative learning for achieving optimal control policies in energy systems, advancing the goals of sustainable and efficient smart grids.
keywords:
smart grids, federated learning, reinforcement learning, energy efficiency
journal:
[…]
{Frontmatter}
[…]
Cuadrado et al.
[…]
The concept of smart grids involves a series of strategies that encompass new dynamics that the electricity grid should integrate to adapt to the requirements introduced by several factors: climate change, increased integration on renewables, changes in regulations, and the need to improve energy efficiency, among others Ton and Smith (2012). Each component of that vision brings challenges we need to solve to make it feasible, and we are tackling one of them in the present work: optimal control of energy assets in microgrids, which is essential to achieve zero energy waste and to guarantee fast integration of renewables without affecting the reliability of the supply. This research impacts Sustainable Development Goals 7 and 11.
1 Introduction
Year after year, the majority of countries renew their commitment to the Sustainable Development Goals (SDGs) and the Paris Agreement in 2015 by evaluating the advancement in the established ambitious targets to combat climate change, with the eye set on the objective of limiting global warming to below 2°C, ideally 1.5°C, compared to pre-industrial levels. However, the most recent report from the Intergovernmental Panel on Climate Change (IPCC) IPCC (2023) reveals that humanity’s efforts remain insufficient, with many targets far from being achieved, which has become a constant. Many challenges need urgent action, and we should increase our efforts to get back on track. Machine learning (ML) is a technology that, when used correctly, can accelerate the actions needed to fight global warming. ML can support the transition from fossil fuels, accelerate renewable energy adoption, improve energy efficiency, and promote sustainable transportation, offering powerful tools to meet some of the most pressing climate goals Rolnick et al. (2019).
Integrating renewable energy sources like wind and solar into the energy grid is a significant challenge due to their intermittent and unpredictable nature, complicating real-time grid balancing. Fluctuations in renewable energy production can lead to supply-demand mismatches, necessitating advanced solutions such as energy storage systems and predictive analytics Perera et al. (2020); Gernaat et al. (2021). Smart Grid technologies, which leverage data-driven approaches, offer a way to address these complexities. By monitoring and managing the energy supply chain from generation to consumption, Smart Grids aim to optimize operations for diverse stakeholders, minimize costs, reduce greenhouse gas emissions, and ensure grid stability, reliability, and resilience.
Artificial Intelligence (AI) has gained recognition as a powerful tool for addressing the urgent threats of climate change. Its ability to process and analyze vast datasets allows for better comprehending environmental shifts and developing strategic solutions. Reinforcement Learning (RL), a subset of AI, offers some capabilities that are especially relevant to addressing the complexities of energy systems. Namely, its ability to handle non-linear, dynamic, and uncertain systems makes it particularly effective for managing distributed energy resources and microgrids. Model-free RL algorithms learn from interactions between an agent and its environment, enabling them to identify optimal control actions even under dynamic environments with constantly changing conditions without depending on a precise model, which is the traditional approach.
However, RL also faces challenges that hinder its applicability in real life. It is a method that relies heavily on massive sampling, making a data-hungry approach; as a consequence of its sample inefficiency, its performance decays when used in out-of-the-distribution (OOD) data (generalization) and particularly for applications in the energy efficiency field, RL needs to access sensitive data on energy usage to reach optimal solutions, creating the need of implementing privacy-preserving methodologies with a minor impact on the utility.
This work studies the application of RL within microgrids, which is one of the strategies that belong to the smart grid concept. The Department of Energy of the United States provided a commonly accepted definition for microgrid: “a group of interconnected loads and distributed energy resources within clearly defined electrical boundaries that act as a single controllable entity concerning the grid. A microgrid can connect and disconnect from the grid to operate in grid-connected or island mode” Ton and Smith (2012). For this purpose, we defined a custom microgrid scenario using a well-known environment called CityLearn Vázquez-Canteli et al. (2019), which allowed us to study the properties and challenges of RL and how collaborative learning can improve the learning process.
We propose a method based on two commonly used methodologies in the ML domain: RL and federated learning (FL). The first follows the intuition that model-free learning is a feasible alternative to managing distributed energy resources (DERs) in a microgrid. More specifically, we picked the Trust Region Policy Optimization (TRPO) Schulman et al. (2015) algorithm to enjoy the benefits of its adaptive step size, considering we are not using large models. The second is to profit from having multiple agents learn in parallel while preserving their privacy. This framework, which we named FedTRPO, harnesses inherent generalization capability and privacy guarantees from FL to discern latent relationships within the feature space. We present a summary of our contributions next:
  • We present a customization to the CityLearn environment that allows a better understanding of the learned policies and the impact of the reward function in this case. We also modified some characteristics that better represent the dynamics of real-life scenarios.
  • This work introduces the FedTRPO learning framework and evaluates it in the context of our proposed smart grid problem, leveraging the faster convergence of TRPO with the fewer communication rounds time of FL.
  • Our results demonstrate the generalization capabilities of FL and how reasonable considerations like personalization and split-learning Han et al. (2022) can reduce the optimality gap in the learned policies.
  • We compared our framework against state-of-the-art RL methodologies in a standard FL setup.
  • We provide a wrapper for the CityLearn environment that makes it compatible with TorchRL Bou et al. (2023), improving the modularity and portability of our code. We also provided a version of the TRPO algorithm using TorchRL.
2 Related Work
Federated Learning. FL provides a collaborative learning framework relying on clients that share the same model architecture but learn locally based on their portion of the data distribution. After some training steps, reducing the clients’ models generates a new iteration of the original model that combines the individual learnings. Federated Averaging (FedAvg) McMahan et al. (2017) stands as the pioneering framework in FL, which averages the weights using a centralized server, often called the parameter server (PS). Other approaches offer a decentralized way of averaging the models; for instance, the authors of Koloskova et al. (2019) introduced a methodology that consists of direct communication between clients for several rounds until they all update their model with the reduced one, which they called Gossip algorithms. Further research on FL tackles some of the challenges of FedAvg, like the suboptimal nature of averaging models under heterogeneity of the clients (difference in the portion of the data distribution they access) or the privacy loopholes created by FedAvg, contrary to the belief that it preserves the privacy of the clients. The work in Vepakomma et al. (2018) presents an interesting approach named split-learning, which tackles personalization (adjusting the combined model to the client’s characteristics) and privacy by keeping a local private part of the model not shared for aggregation. Differential Privacy (DP) Geyer et al. (2017) is another technique often used with FL. It consists of offuscating the original data by passing it through a noise mechanism that ensures a level of utility while ensuring the privacy preservation of individuals or groups to a certain extent. Other topics of interest in the domain are learning in heterogeneous setups Diao et al. (2021) and analysis of the generalization capabilities of FL Yuan et al. (2022). With the proper considerations, FL can handle heterogeneous data training, preserve privacy, and generalize on unseen data, making it a promising methodology for addressing challenges related to smart grid scenarios.
RL-based and Federated RL-based microgrids. The microgrid scenario increasingly receives attention in research on the intersection of ML and energy. The most common setup consists of a group of homes constituting a microgrid. Each one of them has its load profile (which could be controllable or not), energy generation resources (generally solar panels or wind turbines), and storage devices (electricity, heating, or cooling.) There are community efforts that enable tackling this scenario, and the CityLearn environment Vázquez-Canteli et al. (2019); Nweye et al. (2024) is the most recurrently used. It provides an interface compatible with Farama Foundation’s Gymnasium Towers et al. for building energy coordination, vehicle-to-grid control, thermal control, and demand response based on a real dataset Wilson et al. (2021). Several contributions used it in the research of RL; for instance, the work presented in Nweye et al. (2023) explored a multi-agent RL approach for training agents to control the battery energy storage systems (BESS) pursuing the minimization of a weighted sum of a set of metrics of interest for energy user and grid operator: electricity price, carbon emissions, zero net energy, average daily peak, and load factor. To do so, they implemented a multi-agent version of the Soft Actor-Critic algorithm with a shared replay buffer. They employed three different setups for the learning case, differing in the number of agents contributing to the memory buffer and the data points used for the training. The authors of Nweye et al. (2022a) explored the challenges of multi-agent RL using the same environment, showing that a warm start with rule-based control (RBC) policies seems to be the best way of directing the RL learning to find feasible RL policies. However, other contributions leveraged their environments using private datasets of energy usage, like the work presented by Ceusters et al. (2021), which adopted a model-based reinforcement learning (RL) approach tailored to multi-energy systems, enhancing it with constrained optimization algorithms to derive optimal solutions. Meanwhile, Su et al. (2022) introduced a Federated RL (FRL) methodology, considering the dynamics between Energy Data Owners (EDOs) and Energy Service Providers (ESPs), addressing free-rider phenomena and non-independent and identically distributed (IID) data distributions. Furthermore, Cuadrado et al. (2023a) introduced a hierarchical FRL, which demonstrated advantages in the learning process while integrating the FedAvg method in an existing hierarchical RL setup detailed in Cuadrado et al. (2023b).
Although we consider that the existing work has been paving the foundations of RL applicability in real life, we find two relevant concerns. First, many contributions aim to see a monotonic increase behavior of the return function (the main objective of RL) to claim that a policy is optimal when that is not necessarily the case. One of the main challenges of this scenario is avoiding the local minima from trivial cases (e.g., not doing anything, always charging, constantly discharging, on-off policies), and an increased return could not mean optimality, not even gradients of magnitude close to zero mean optimality. Second, we noticed that in the contests organized by the same authors of CityLearn in the last years, the winning teams disregarded the impact of the emissions, prioritizing other metrics that had higher weight in the final score function Nweye et al. (2022b); Nagy et al. (2021). We found that in the context of climate change, the reduction of emissions should be a priority, and we might be ignoring why this scenario seems harder to resolve when prioritizing emission reduction.
3 Method
3.1 Environment
As mentioned before, we used CityLearn environment to try our framework, however, instead of using the standard data they include in their package we decided to generate synthetic data that allow us to understand clearly the learned policies through simple cases. For this work we focused on data representing homes that can be self-sustainable, in other words, each home can provide enough energy to serve its demand, so the optimal solution must have zero cost and emissions. Our data generation process takes as a reference the citylearn_challenge_2022_phase_all scenario. It modifies the time series for non-shiftable load to ensure solar energy can cover the demand profile completely. We included other modifications to the environment:
  • We added a new logic for the selling energy rate. We found it unrealistic that in the original scenario, the users receive a compensation equal to the Time-of-Use (ToU) tariff. This setup discourages houses from being self-sustainable, as we will detail later.
  • We modified the capacity of the BESS devices. We simplified their models so they don’t suffer degradation or decay in their performance, helping to ease the understanding of the policies we are training, which is one of the main interests of this research.
  • We introduced additional observations representing load predictions for some steps ahead, which we consider essential to finding optimal policies.
[…]
Figure 1: Optimal Actions and State of Charge: Mean and variance throughout the day, considering a complete year, for a building. By design, we know the optimal policy that must yield zero emissions and cost beforehand.
We kept the grid price pattern, which fluctuates throughout the day, peaking during high-demand hours, while the solar panels contribute energy during daylight hours. Allowing the agent to stockpile energy during off-peak times and, when solar generation exceeds immediate needs, later release stored energy during high-price periods. Emissions follow the actual emission rate from the grid mix (kgCO2e/kWh) for the dataset used in the environment. With our process, we can generate training data up to the number of buildings existing data in the reference case. We generate convex combinations of the training buildings’ time series for the validation data. Each episode spans twenty-four hours, and $t$ is the index for the time step. For a building $i$, an agent considers the accessible observations to decide the charge or discharge of the battery in a time step ($E^{\text{batt}}_{t,i}$), aiming to meet the building’s load requirements. Its primary objective is minimizing its house’s overall energy cost and emissions and, as an extension, for the whole microgrid. The table 3.1 presents an overview of the RL case.
[…]
Table 1: RL case for agent $\mathbf{i}$: detail of the state composition with the units of each feature, action definition, and reward design.
State: $\mathbf{s_{t,i}}$
Hour $t$$\in~{}\{1,2,\ldots,23\}$-
[…]
Other Relevant Values
Battery Capacity $B_{i}^{\text{cap}}$ $\in~{}\mathbb{R}_{\geq 0}$ $kWh$
[…]
Action: $\mathbf{a_{t,i}}$
[…]
Reward: $\mathbf{r_{t,i}}$
$P_{i,t}=B_{i}^{\text{cap}}*\begin{cases}\max\left\{0,B^{\text{action}}_{t-1,i}-(1-B_{t-1,i}^{\text{soc}})\right\}&~{}~{}\text{if}~{}~{}B^{\text{action}}_{t-1,i}\geq 0\\ \max\left\{0,|B^{\text{action}}_{t-1,i}|-B_{t-1,i}^{\text{soc}}\right\}&~{}~{}\text{o.w.}\end{cases}$
[…]
Our design of the reward function is a weighted sum of the resulting cost and emissions for the complete episode (a full day), and we established the values as $W_{C}=0.4$ and $W_{G}=0.6$ to enforce the prioritization of emission reduction. Previous works, namely Nweye et al. (2023), opted to make this weighting a tuned hyperparameter, which ended by defining $W_{G}=0$, giving full priority to energy cost. Still, the authors claimed that their learned policy reduced emissions when just caring about minimizing energy costs. We found that the claim might not be accurate, as in their scenario, the electricity pricing and the selling pricing were the same, promoting that the agents buy energy from the grid to charge the battery to sell it, then when the price is higher, generally at night.
Inspired by the same work, we added a penalization to the reward function as well, but ours aims to guide the agent to avoid unfeasible actions; as we are going to present later, we parameterized the continuous action using a normal distribution, which could fall outside the range of the desired actions. We also found that the penalization used in Nweye et al. (2023) would lead to suboptimal policies as it promotes the use of the battery in every step, even in scenarios where the agent should save energy for future usage when more convenient.
3.2 Synthetic Data
For this research, we decided to generate our data as it allowed us to explain better the policies we learned. Although using accurate data is always desirable for ML, there is low availability, especially regarding energy usage patterns, which concerns privacy. For example, the CityLearn environment constructs its 2022 scenario using data from 17 Sierra Crest Zero Net Energy community buildings from another study about ToU strategies Narayanamurthy et al. (2016). They couldn’t access all the accurate data details, so they had to generate data that emulated the behavior based on some prior information about the ToU strategies the households might have used. Research about microgrid control systems is not particularly affected by lack of data, as prior information about consumption behaviors, typical weather conditions given a location, seasonal changes, and factors influencing energy use are accessible. Synthetic data could represent typical scenarios an agent may face and have the advantage of creating relevant evaluation scenarios by adding noise or making a convex combination of seen data.
[…]
(b) Normal Dataset
Figure 2: Sample of the Mean and Variance of Solar Generation and Non-Shiftable Load over 24 hours for a year of data for one of the buildings.
We generated our synthetic data by altering the Non-Shiftable Load time series using a simple rule that defines the policy we want to learn: the households have enough energy and battery capacity to serve their demand when there is solar energy and to store what is necessary when there is no sun. By definition, our scenario leads to net-zero energy consumption. Figure 3.1 shows how two versions of the data we generated, in the case of LABEL:fig:simple_solar_load, we simplified the solar profile by making it constant through the sunny hours. For 2(b), we kept the solar profile as in the original reference case from CityLearn. By simplifying these profiles, we intended to reduce the state space for the policy while still keeping it challenging to learn. Another critical difference in our scenario is the pricing for energy. Our reference scenario from the CityLearn environment used the real pricing time series used in the Zero Crest Community Study. Still, for the environment implementation, they defined that the selling price is the same tariff whenever a household sends back energy to the grid. By inspecting 3(a), it is easy to notice that a policy optimizing exclusively by the energy cost would charge as much as possible in the low-price hours to sell energy when it is almost twice the buying price. In our scenario, we defined a margin of 40% of the lowest daily electricity pricing to ensure we don’t learn policies that bypass the weighting factor $W_{G}$ by finding profitable energy trades.
[…]
(a) Original scenario
[…]
(b) Our scenario
Figure 3: Mean and Variance of Electricity Pricing (ToU) and the Selling Pricing over 24 hours for a year of data.
3.3 Reinforcement Learning
The RL component of our methods happens at the local training. We compared TRPO against a state-of-the-art algorithm called Proximal Policy Optimization (PPO) Schulman et al. (2017), which emerged as an alternative to TRPO that avoids the use of second-order information. PPO relies on the same theoretical base as TRPO but redefines the policy optimization objective as an unconstrained expression. The scenario we study in this work, and the case of microgrids’ agents, does not require the use of very deep neural networks, and that is a reason why we believe that second-order methods are feasible in these cases, especially for RL methods, which tend to be very sensitive to the learning rate. TRPO explores the vast space by identifying suitable step sizes within a confined region, ensuring stable policy, monotonic improvement, and more effective exploration than other policy gradient methods.
Another essential aspect of opting for an RL approach is defining a proper reward function, especially considering that we defined a continuous action space, which means the number of possible actions is unbounded. The typical approach for this case is using a policy network that outputs the parameters of a Gaussian distribution (re-parameterization trick) to sample a real value that defines the action from that distribution. Throughout the training process, we don’t have guarantees that the policy network will explore feasible actions (within the range specified in 3.1), especially in the first iterations. That fact motivated the addition of a penalization $P_{i,t}$ that impacts the return of a trajectory (the accumulated sum or rewards.) An essential consideration we made when designing our reward function is that we are not guiding it to follow a specific policy we know beforehand; we are just guiding the exploration of the agent by avoiding ineffective actions.
3.4 Models
The present work explores the advantages of a collaborative learning approach for the case of microgrids, given their distributed nature: each household has an agent that explores and learns from different sections of the state space given the individual attributes of the household (generation, load profile, battery capacity). For the present work, we have two types of models for both Policy and Critic networks: a shallow neural network and a custom model that adds a grouping feature logic as the first part of the feature extraction; we refer to it as Personal Encoding in Figure 4.
[…]
Figure 4: Overview of our models: We use the same type of network for Policy and State-Value networks, changing only the output layer. The Personal Encoding block is optional.
Each household’s agent could have its policy network or share a common one based on the desired configuration, and we defined a single critic network for all of them as we need it only during training. The personal encoding block represents a part of the network that remains local; the FedAvg process does not access it. It accesses features that should remain private (e.g., number of people living in the household, ToU strategy, appliances, area of the house, etc.) and learns a latent representation, used as an input to the public model, stacked with the public features, which aligns with the concept of split-learning Vepakomma et al. (2018), which we can assume as a personalization strategy in this use case; each house has an adaptor that allows the policy to perform better in its case.
3.5 Federated RL
An actual FL setup implies having clients interact through a communication network channel throughout the training rounds. In the context of our study case, we can imagine it as the agents operating in an embedded device that controls the batteries of each household and connects to a parameter server through a dedicated network connection. However, this case would introduce some of the challenges of FL: limited bandwidth, faulty nodes or updates, byzantine clients, etc. The main interest of this work doesn’t aim to alleviate those challenges but to explore the benefits of collaborative learning. Thus, we are not required to implement the exact flow of FL to perform our experiments.
As Figure 4 shows, we stack the features (observations and personal encoding) for all the agents and pass them to their corresponding networks. Passing the stacked observations to a single network has the same effect as doing a FedAvg: each agent samples their trajectories, and after a forward pass, the backward pass will propagate the average gradient of all of them, which is, in practice, the same behavior we expect from the FL methodology. We use a simple one-hot encoder for each home as the personal encoding. Although we are not in the process of training a private part of the policy network, we can explore the exact effects of personalization by having a vector representation that identifies each home. Using FL is to discern the interconnected relationships in the observations across buildings collectively. Relying solely on data from a single building risks overfitting, but leveraging diverse data distributions from multiple buildings helps uncover latent connections. Moreover, the model adeptly extrapolates unseen data instances by assimilating insights from disparate building data distributions, handling anomalies emerging from using usage patterns.
4 Experiments
4.1 Preliminary Results
The present paper is an extension of the work in Li et al. (2024), in which we found results that suggested the generalization ability of our FedTRPO approach by employing other synthetically generated training and testing datasets. We defined our baseline, denoted as Upperbound, by overfitting our testing data. Then we tried these scenarios: training individual agents for each building denoted as Ind. Agent and training with FedTRPO with and without personal encoding. Figure 5 presents five buildings’ average reward and emission trends.
[…]
Figure 5: Previous Results: Average reward and emission of the five buildings across five random seeds. Upperboud (Blue): A single TRPO agent trained using the testing dataset to establish the upper-performance limit. FL (Green): Model structured with all parts shared trained with FL methodology. Ind. Agent (Red): TRPO agent trained separately for each building. FL Personalization (Orange): FedTRPO with personalized encoding as detailed in Section 3, trained using FL methodology.
Our interpretation at the time suggested that in Figures 5, FedTRPO consistently outperforms FL and Ind. Agents across all buildings, approaching the baselines we believed optimal. However, further evaluations of the policy revealed that our upper bound was still far from optimality. It was a local minimum close to a naive policy: not using the battery. Motivated by this, we decided to explore the generalization of our FedTRPO approach with more straightforward but more explainable scenarios.
4.2 Simple Policy Learning
The main set of experiments of the current work aims to test the capability of the same approach we explored in the preliminary work in a simplified scenario that remains challenging and relevant: net-zero policy learning, detailed in 3.1. By doing this, we can refine components of our original approach, like the reward function and the improvements to the environment we mentioned before, as the original experiments used the same setup as in the out-of-the-box CityLearn environment.
4.2.1 Experiment Setup
We generated two environments for our experiments using the simplified version of the data LABEL:fig:simple_solar_load: a microgrid with two houses and one with five houses. Each one also has a shifted version that creates day gaps between the house’s time series, simulating that each house is subject to different environmental conditions and solar irradiance patterns. Regarding the model configurations, as described 4, we run experiments in four setups: shallow network, shallow network with personal encoding, shallow network with grouping, and shallow network with personal encoding and grouping. We defined four group features:
  • Personal encoding with load-related observations: $E_{t,i}^{\text{load}}$.
  • Personal encoding with solar-related observations: $E_{t,i}^{\text{solar}}$, $I_{t,i}^{\text{direct}}$.
  • Pricing with predicted observations: $R_{t,i}^{\text{buy}}$, $R_{t,i}^{\text{sell}}$, $E_{t,i}^{\text{load+4}}$, $E_{t,i}^{\text{load+6}}$, $E_{t,i}^{\text{load+12}}$, $E_{t,i}^{\text{load+24}}$, $I_{t,i}^{\text{direct+4}}$, $I_{t,i}^{\text{direct+6}}$, $I_{t,i}^{\text{direct+12}}$, $I_{t,i}^{\text{direct+24}}$.
  • The rest: $t$, $d$, $B_{t,i}^{\text{soc}}$, $E_{t,i}^{\text{net}}$, $R_{t,i}^{\text{carbon}}$.
We perform a hyperparameter search with Wandb sweeps Biewald (2020), using Bayesian search to find the best configurations for PPO and TRPO with the following final configurations:
Table 2: Final training hyperparameters for both RL methodologies. Parameters considered in the sweep have a *.
[…]
(a) PPO
Hyperparameter Value
Communication rounds $30$
[…]
(b) TRPO
Hyperparameter Value
Communication rounds $30$
[…]
For our comparison, we fixed the communication rounds for both algorithms to evaluate how close they can come to the optimal solution within the same number of iterations. Another essential detail is setting the discount factor $\gamma=1$, meaning that our agents don’t consider a discounted reward, given that we are running episodes of just one day (24-time steps). That is an important decision we considered an improvement concerning previous work in similar scenarios: in Nweye et al. (2022a, 2023), they trained episodes of a complete year while selecting $\gamma<1$, which suggests that agents were short-sighted and that long trajectories are not informative in estimating the value functions. About TRPO, the only tunable parameter is the critic’s learning rate, but we took it from the PPO process. The rest of the configurations followed theoretical insights. Regarding the hyperparameter $\lambda$, we also set it to $1$ to avoid bias towards a specific advantage estimation in a given timestep.
4.3 Results Summary
When analyzing the results, it is essential to consider that the PPO approach was tuned and used the best set of hyperparameters we found, while TRPO was not tuned. Both techniques rely on the same theoretical base: both fall under the policy gradient-based methods of RL, both aim to optimize a surrogate function for the return based on the Performance Difference Lemma (PDL) proposed in Kakade and Langford (2002), both aim to do conservative updates to the policy parameters to ensure a stable learning process using the Kullback-Leibler (KL) Divergence between the iterations of the policy’s distribution. The main difference lies in removing the use of second order information in the optimization process (TRPO), replacing it with a clipped version (PPO). Each approach has its advantages and disadvantages. We faced them during the training process and also reflected in the results. For instance, PPO faced challenges when trying to reduce the optimality gap in the vicinity of the solution, and adding more iterations destabilized the training (even using a small learning rate). TRPO suffered when facing a nonconvenient initialization, which is a common problem in second-order methods, limiting its capability of exploiting sampling (more trajectories would lead to a better Hessian estimation) to optimize in fewer iterations with an adaptive learning rate.
4.3.1 Quantitative Results
The tables 3 and 4 show the experiments we mentioned in the previous section: base means we used the shallow network, pe that we included personal encoding, and gf that we grouped features. We ran each combination for five different seeds. We averaged the results, making $160$ experiments with an average run time of $70m$ for the PPO runs, and $110m$ for the TRPO runs using an NVIDIA RTX 4000 Ada Generation GPU.
Starting with the environment containing just two buildings, we found that training with the shifted variant is more challenging than the normal one, which makes sense considering we provide more diverse data without increasing the minibatch size. We cannot conclude from the results that using personal encoding or grouped features made a difference in the case of PPO, but for TRPO, it seems that they could be beneficial, as was evident in the shifted dataset. In general, as expected, the performance of the tuned PPO was superior to that of the TRPO, and the reward function value demonstrated it.
[…]
Table 3: Experimental Results for 2-Building Environment: Values in the table represent the average for $5$ seeds. Values in parentheses represent the difference concerning not having battery management.
Experiment Train Cost Eval Cost Train Emissions Eval Emissions Train Reward Eval Reward
PPO base 0.04227 (-0.08977) 0.04099 (-0.09148) 0.06380 (-0.07487) 0.06260 (-0.07652) -0.19331 -0.18754
gf 0.03054 (-0.10150) 0.03049 (-0.10199) 0.04341 (-0.09526) 0.04308 (-0.09604) -0.13227 -0.13188
pe 0.03885 (-0.09319) 0.04630 (-0.08618) 0.05841 (-0.08026) 0.06384 (-0.07527) -0.17686 -0.19956
pe gf 0.03051 (-0.10153) 0.03227 (-0.10021) 0.04306 (-0.09561) 0.04637 (-0.09274) -0.13065 -0.14021
[…]
gf shifted 0.24691 (+0.10051) 0.25710 (+0.11419) 0.28074 (+0.08558) 0.30601 (+0.10943) -1.46458 -1.51916
pe shifted 0.04541 (-0.10099) 0.10654 (-0.03637) 0.08211 (-0.11305) 0.13600 (-0.06058) -0.21291 -0.56921
pe gf shifted 0.10324 (-0.04316) 0.15339 (+0.01049) 0.10922 (-0.08594) 0.17679 (-0.01979) -1.10813 -1.34919
TRPO base 0.31658 (+0.18454) 0.32391 (-0.19143) 0.26502 (+0.12635) 0.26908 (+0.12997) -1.12509 -1.12524
gf 0.31882 (+0.18678) 0.32288 (+0.19040) 0.27001 (+0.13134) 0.27020 (+0.13108) -1.02230 -1.01532
pe 0.10567 (-0.02637) 0.10402 (-0.02846) 0.11511 (-0.02357) 0.11271 (-0.02641) -0.47736 -0.47064
pe gf 0.10519 (-0.02685) 0.10355 (-0.02893) 0.11519 (-0.02348) 0.11302 (-0.02609) -0.47877 -0.47164
[…]
gf shifted 0.17848 ( 0.03208) 0.16951 (+0.02660) 0.19592 (+0.00076) 0.18621 (-0.01037) -0.77824 -0.74692
pe shifted 0.14833 (+0.00193) 0.14171 (-0.00120) 0.17103 (-0.02413) 0.16301 (-0.03357) -0.64225 -0.63830
pe gf shifted 0.12728 (-0.01912) 0.12032 (-0.02259) 0.15287 (-0.04229) 0.14513 (-0.05145) -0.57360 -0.56441
In the case of the environment with five buildings, grouping features seems to play a more critical role than in the case of just having two buildings. We interpret this as the effect of having more agents, which means more sampling and more diverse data, which can help the networks unveil patterns in the data that are not evident by simple inspection. However, this does not seem to be the case for TRPO, in which the simple model yields the best performance, although it is essential to note that the performance of TRPO is poor based on the reward function.
[…]
Table 4: Experimental Results for 5-Building Environment: Values in the table represent the average for $5$. Values in parentheses represent the difference concerning not having battery management.
Experiment Train Cost Eval Cost Train Emissions Eval Emissions Train Reward Eval Reward
PPO base 0.04001 (-0.09140) 0.03856 (-0.09297) 0.06094 (-0.07710) 0.05949 (-0.07866) -0.18392 -0.17930
gf 0.24239 (+0.11098) 0.25020 (+0.11867) 0.23766 (+0.09962) 0.24276 (+0.10461) -1.55780 -1.55296
pe 0.04089 (-0.09052) 0.04505 (-0.08648) 0.06159 (-0.07645) 0.06488 (-0.07328) -0.18951 -0.20773
pe gf 0.02889 (-0.10253) 0.03233 (-0.09920) 0.04235 (-0.09569) 0.04711 (-0.09104) -0.13100 -0.14662
[…]
gf shifted 0.03725 (-0.11299) 0.05633 (-0.09414) 0.05421 (-0.10924) 0.07025 (-0.09474) -0.16773 -0.22911
pe shifted 0.06305 (-0.08719) 0.08847 (-0.06200) 0.09080 (-0.07265) 0.10826 (-0.05674) -0.28214 -0.39795
pe gf shifted 0.10324 (-0.04316) 0.15339 (+0.01049) 0.10922 (-0.08594) 0.17679 (-0.01979) -1.10813 -1.34919
TRPO base 0.09811 (-0.03330) 0.09737 (-0.03416) 0.11173 (-0.02630) 0.11086 (-0.02730) -0.46143 -0.45768
gf 0.10437 (-0.02704) 0.10312 (-0.02841) 0.11455 (-0.02349) 0.11327 (-0.02488) -0.47631 -0.47153
pe 0.09843 (-0.03298) 0.09750 (-0.03403) 0.11019 (-0.02785) 0.10906 (-0.02909) -0.46143 -0.45614
pe gf 0.10177 (-0.02964) 0.10044 (-0.03109) 0.11274 (-0.02529) 0.11140 (-0.02676) -0.46460 -0.45958
[…]
gf shifted 0.17795 (+0.02771) 0.16457 (+0.01410) 0.16254 (-0.00091) 0.15295 (-0.01204) -0.77981 -0.74353
pe shifted 0.12846 (-0.02178) 0.11979 (-0.03068) 0.13076 (-0.03269) 0.12391 (-0.04109) -0.56240 -0.55708
pe gf shifted 0.14833 (+0.00193) 0.14171 (-0.00120) 0.17103 (-0.02413) 0.16301 (-0.03357) -0.64225 -0.63830
PPO overperforms TRPO significantly, but it is essential to highlight that we didn’t spend time fine-tuning it. Still, it managed to identify some patterns in the data, as we will show in the qualitative results.
4.3.2 Quantitative Results
We inspected the best-performing policies, considering the eval. reward, according to Tables 3 and 4 to understand if we accomplished the objective of beating the baseline policy, which is not using a battery. In the case of PPO, we can notice in 6 that it was the case. However, we can see an effect we mentioned before in Figures 6(a) and 6(b): extending the training to reduce the optimality gap destabilizes the training, which is something that happens to gradient-based optimizers with a fixed learning rate.
[…]
(a) 2B Eval. Energy Cost
[…]
(b) 2B Eval. Emissions
[…]
(c) 5B Eval. Energy Cost
[…]
(d) 5B Eval. Emissions
Figure 6: Best PPO policies: Plots show mean with an envelope of one standard deviation computed from the five different seeds.
Increasing the number of buildings makes it harder for the agents to approach net-zero energy and training with the shifted version of the dataset, something expected. A first inspection of the results for the five-building environment might give the impression that the training process was more stable, but the reason must be that the policy didn’t get as close to net zero as in the case of two buildings; the instability in the training happens in the vicinity of the optimal solution.
[…]
(a) 2B Eval. Energy Cost
[…]
(b) 2B Eval. Emissions
[…]
(c) 5B Eval. Energy Cost
[…]
(d) 5B Eval. Emissions
Figure 7: Best TRPO policies: Plots show mean with an envelope of one standard deviation computed from the five different seeds.
For the best-performing policies using TRPO, we can notice how the low returns reflect high variance among runs. However, some configurations (model with personal encoding and grouping) limited the variance. They beat the baseline on each run, which is a highlight considering the lack of tuning. That event was present in both environments.
[…]
(a) PPO: B0 Policy
[…]
(b) PPO: B0 SoC
[…]
(c) PPO: B1 Policy
[…]
(e) TRPO: B0 Policy
[…]
(f) TRPO: B0 SoC
[…]
(g) TRPO: B1 Policy
[…]
(h) TRPO: B1 SoC
Figure 8: Best policies for the 2-buildings environment: Comparison of learned policy against the optimal policy, averaged for the five seeds.
In Figure 3.1, we notice that PPO gets very close to the optimal policy, but a few misalignments explain why it couldn’t reach net zero. From the best TRPO policy, we can highlight that although it’s far from optimal, it is evident that it learned a pattern and that although sub-optimal, it’s still a policy that beats the baseline.
5 Conclusion
The present work presents a systematic study of the learning capabilities of federated RL in the context of distributed control systems for microgrids, particularly in the case of batteries. We introduced a simple but expressive evaluation framework based on a simplified scenario that, by design, has an optimal policy that yields net-zero: no emissions and no cost. In other words, a self-sustainable microgrid. We introduced FedTRPO, a federated reinforcement learning framework that combines Trust Region Policy Optimization (TRPO), FedAvg, and personalization through split learning to optimize energy management in our case. Our framework addresses key challenges in distributed energy systems, including privacy preservation, policy optimization for heterogeneous energy profiles, and achieving net-zero energy consumption.
Our experimental results show that FedTRPO performs competitively compared to the state-of-the-art PPO operating in a federated setup. While PPO demonstrated superior performance in some instances, TRPO provided improved convergence without hyperparameter tuning. We believe the best solution would involve a mix of the two RL algorithms to address their challenges. PPO faces difficulties when it is close to the optimal solution due to the fixed learning rate, and a scheduler might be a solution for some instances, but it becomes a new hyperparameter. TRPO relies entirely on sampling to converge in a few steps to the optimal. Still, the quality of the first trajectories and the initialization of the parameters limit its best performance, which is a common problem of second-order optimization methods. The simple solution could be performing a few iterations of PPO to create a better initialization point and then continue with TRPO, profiting from its adaptive step size, which will play an essential role in the vicinity of the optimal solution.
This study demonstrates the feasibility and effectiveness of federated reinforcement learning for smart energy systems, particularly in scenarios requiring both performance and privacy. We identified that learning simple optimal policies is more challenging than expected, but learning simple cases allows for better refining of the methodologies, which could positively impact the learning process with actual data. More importantly, simple environments could be the base of a curriculum-based approach to learning more complex cases in real-life scenarios. By creating a set of designed cases that represent desirable behaviors of the policies, we could make a more stable process to enable RL in this kind of case, which seems to be a challenge previous work in the domain faced: learning a purely RL-based policy to manage distributed energy resources is extremely hard.
{Backmatter}
Acknowledgments
We thank Yunxiang Li for their contributions to the previous work, which was a foundational basis for the present study. We want to congratulate Vassilis Choutas for his implementation of the Trust-Region optimization method, accessible here, which was key to implementing our TRPO version.
Funding Statement
This work did not receive funding from any specific grant, whether from governmental agencies, commercial entities, or non-profit organizations.
Competing Interests
The authors declare none.
Data Availability Statement
We utilized synthetically generated data, accompanied by Jupyter notebooks that enable others to reproduce the data creation process. Additionally, all the code, including the custom modules explicitly developed for this paper, has been made publicly available at the following repo, allowing for the replication of experiments.
Ethical Standards
The research meets all ethical guidelines, including adherence to the legal requirements of the study country.
Author Contributions
Conceptualization: N.C.; M.T. Methodology: N.C. ;S.H ;M.T. Data curation: N.C.; M.T Data visualisation: N.C. Writing original draft: N.C. All authors approved the final submitted draft.